40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 26 02:15:36 2020, gautam, Update, LSC, Lock acquisition portal entry 
    Reply  Tue May 26 02:31:00 2020, gautam, Update, LSC, Lock acquisition sequence PRFPMIlock_1274418200_1274418550.pdf
       Reply  Wed May 27 20:14:51 2020, Koji, Update, LSC, Lock acquisition sequence 
    Reply  Tue May 26 02:37:19 2020, gautam, Update, LSC, DARM loop measurement and fitting DARM_TF.pdfDARM_TF_breakdown.pdf
    Reply  Tue May 26 03:01:35 2020, gautam, Update, LSC, CARM loop  CARM_OLTF.pdf
       Reply  Wed Jun 3 01:46:14 2020, gautam, Update, LSC, CARM loop  CM_loop_topology.pdfCARM_TFs.pdfCARM_OLTF.pdfCARM_xover.pdfCARM_OLG_evolution.pdf
    Reply  Tue May 26 03:06:59 2020, gautam, Update, LSC, PRFPMI sensing matrix PRFPMI_20200524sensMat.pdfPRFPMI_20200524sensMatHistograms.pdf
    Reply  Tue May 26 03:26:58 2020, gautam, Update, LSC, Preliminary noise budget PRFPMI_NB.pdf
    Reply  Tue May 26 14:32:44 2020, gautam, Update, LSC, Arm transmission RIN armRIN.pdf
       Reply  Wed May 27 19:36:33 2020, Koji, Update, LSC, Arm transmission RIN 
          Reply  Thu May 28 18:36:45 2020, gautam, Update, LSC, Arm transmission RIN PRFPMIcorner_ASC_PIT_1274419354_1274419654.pdfPRFPMIcorner_ASC_YAW_1274419354_1274419654.pdfPRFPMIcorner_ASC_coherence_1274419354_1274419654.pdf
             Reply  Fri May 29 00:34:57 2020, rana, Update, LSC, Arm transmission RIN 
             Reply  Wed Jun 3 02:14:32 2020, gautam, Update, ASC, PRC ASC improves arm transmission RIN PRC_ASCsignals.pdfarmRIN_PRC_ASC.pdfPRFPMIcorner_ASC_PIT_1275190251_1275190551.pdfPRFPMIcorner_ASC_YAW_1275190251_1275190551.pdfPRFPMIcorner_ASC_coherence_1275190251_1275190551.pdf
                Reply  Fri Jun 19 16:30:09 2020, gautam, Update, ASC, Some thoughts about ASC sensingResponse.pdfsensingResponse_torque.pdf
                   Reply  Tue Jul 7 14:06:10 2020, gautam, Update, ASC, Some more thoughts about ASC ITM_OL_DCcoupling.png
    Reply  Tue May 26 16:00:06 2020, gautam, Update, LSC, Power buildup diagnostics PRFPMIcorner_DC_1274419354_1274419654.pdfPRFPMIcorner_SB_1274419354_1274419654.pdf
       Reply  Wed May 27 17:41:57 2020, Koji, Update, LSC, Power buildup diagnostics 
          Reply  Wed Jun 3 02:08:00 2020, gautam, Update, LSC, Power buildup diagnostics PRFPMIcorner_DC_1275190251_1275190551.pdfPRFPMIcorner_SB_1275190251_1275190551.pdf
    Reply  Wed Jun 3 01:34:53 2020, gautam, Update, LSC, Lock acquisition update portal 
       Reply  Wed Jun 3 03:29:26 2020, Koji, Update, LSC, Lock acquisition update portal 
          Reply  Wed Jun 3 11:40:56 2020, gautam, Update, LSC, Lock acquisition update portal 
             Reply  Wed Jun 3 18:49:47 2020, gautam, Update, LSC, PRG and CARM signal sign armCavReflectivities.pdfIFOreflectivities.pdfPDHerrSigs.pdfPRGvsLoss_finesse.pdf
Message ID: 15353     Entry time: Tue May 26 03:26:58 2020     In reply to: 15348
Author: gautam 
Type: Update 
Category: LSC 
Subject: Preliminary noise budget 

Summary:

This isn't meant to be a serious budget, mainly it was to force myself to write the code for generating this more easily in the future.

Details:

  • DARM OLTF model from here was used to undo the loop to convert the in-loop measurement to a free-running estimate.
  • The AS55 PD channels were whitened to reduce the effect of ADC noise.
  • To measured channel was 'C1:LSC-DARM_IN1_DQ'.
    • Some care needs to be taken when applying the conversion from counts to meters using the sensing element measured here.
    • This is because the sensing matrix measurement was made using the response in the channel 'C1:LSC-AS55_Q_ERR_DQ'.
    • Between 'C1:LSC-DARM_IN1_DQ' and 'C1:LSC-AS55_Q_ERR_DQ' there is a scalar gain of 1e-4, and a z:p = 20:0 filter.
    • These have to be corrected for when undoing the loop, since the measurement point is 'C1:LSC-DARM_IN1_DQ'. 
  • The "Dark noise" trace was measured with the PSL shutter closed, but all CDS filters up to 'C1:LSC-DARM_IN1_DQ' enabled as they were when the DARM measurement was taken.
  • It would be interesting to see what the budget looks like once the DARM loop gain has been turned down a bit, some low-pass filtering is enabled, and the vertex DoFs are transitioned to 1f control which is hopefully lower noise.
Attachment 1: PRFPMI_NB.pdf  159 kB  Uploaded Tue May 26 04:27:06 2020  | Hide | Hide all
PRFPMI_NB.pdf
ELOG V3.1.3-